Nancy received grades of 72, 83, and 91 on three math exams. What grade must she receive on the fourth exam so that she has at least an 80 average?
Answer by rm29924(97)

let x = the grade Nancy must make on the fourth exam to have an average of at least 80
to get an average you must add the exam grades and divide them by the amount of grades
72+83+91+x/4>=80
246+x/4>=80
multiply both sides by 4 to get rid of it
246+x>=320
subtract 246 from both sides
x>=74
she must receive at least a 74 on her fourth exam